# Detection of entanglement via moments of positive maps

### v1

@@ I Introduction
Another issue in detection of entangled states in experiments is the quantum state tomography with large number of qubits. It is both hard and exhaustive to perform state tomography for many qubits. It was found that partial transpose moments (PT-moments) can quantify quantum correlations and they are more feasible than the state tomography. In addition, it was found that PT-moments are related with randomized measurements and few PT-moments are enough to describe separable states Elben-PRA99 ; Brydges-Sc364 ; Elben-PRL125 ; Huang-NP16 . The local randomized measurements are also proposed to detect multipartite entanglement Ketterer-PRL122 ; Ketterer-Q4 ; Ketterer-PRA106 ; Ohnemus-PRA107 ; Wyderka-PRL131 ; Zhang-arx and bound entanglement Imai-PRL126 . The idea of PT-moments was extended to realignment matrix for a quantum state. It was shown that the moments of a realigned matrix can detect some bound entanglement Zhang-QIP21 . Recently, the relationship between principle minors and PT-moments is worked out where it was shown that all the PT-moments can be written as a simple function of principle minors Ali-QIP22 ; Zhang-AP534 . A more general approach was recently introduced to define the moments of positive maps which are not completely positive Wang-EPJ137 . Based on moments of positive maps, two criteria were proposed to detect entanglement. As transposition is a positive map which is not completely positive, therefore PT-moments criterion is a special case of this general criterion. Although, the criteria based on moments of positive map were able to detect entanglement for a specific family of quantum states, nevertheless there is some range of parameter where these criteria fail to detect entangled states Wang-EPJ137 .
@@
In this work, we propose a slight modification in criterion proposed in Ref.Wang-EPJ137 and show that our suggestion is capable of detecting entanglement completely for several families of quantum states in various dimensions of Hilbert space. We also show that our proposal is capable to detect the complete range of entanglement for a specific family of states where the earlier criterion fails to do so. We find out that reduction map for qubits is equivalent to partial transpose map as the resulting matrices give the same set of eigenvalues for both maps. Consequently, the detection of entanglement is optimal for qubit-qubit system. We also extend the idea of moments of positive map to multipartite systems. In multiqubits systems, reduction map applied to anyone of the qubit is equivalent to partial transpose map. For GHZ and W states mixed with white noise, we find that all the resulting moments for reduction map are exactly same as PT-moments.
@@ II Positive maps and their moments
Let $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ be the bipartite finite dimensional Hilbert space. A mixed bipartite quantum state $\varrho\in\mathcal{H}$ is a density matrix with positive semidefinitive eigenvalues and unit trace. A density matrix is said to be separable (nonentangled) if it can be written as
@@
where $p_{i}\geq 0$ and $\sum_{i}\,p_{i}=1$. In general, it is not easy to verify for a given quantum state whether it is a convex combination of product states. To find out whether a given quantum state is entangled or not, is still an open question despite tremendous efforts in this area of research Horodecki-RMP81 ; Guehne-PR474 . A necessary condition for separability is to check the partial transpose of the density matrix. If $(\varrho^{AB})^{T_{B}}$ is negative (having atleast one negative eigenvalue) then state $\varrho^{AB}$ is entangled. This condition is a necessary and sufficient for $2\otimes 2$ and $2\otimes 3$ quantum systems. However, for higher dimensions of Hilbert space, there are quantum states having positive partial transpose (PPT) nevertheless entangled. Recently, partial transpose criterion was reformulated via partial transpose moments (PT-moments). This was required to address the issue of state tomography with large number of qubits or higher dimensional systems. The PT-moments are defined as
@@
where $k=1,2,\ldots,d$ with $d$ as dimension of Hilbert space $\mathcal{H}$. It was found that we do not have to compute all ${}^{\prime}d^{\prime}$ PT-moment and few moments are sufficient to determine the separability properties of the state. Let $B_{k}(P)$ are $(k+1)\times(k+1)$ Hankel matrices, defined as
@@
A criterion was formulated based on the Hankel matrices that for separable states, Hankel matrices are positive, i.e., $B_{k}\geq 0$. In our recent study, we have observed that for majority of quantum states, only $B_{1}$ and $B_{2}$ are sufficient to determine the separability of a given state. In addition, it is obvious from formulation of PT-moments that they can detect only the NPT region of quantum states and can not detect PPT-entangled states or even genuine entanglement in multipartite settings Ali-QIP22 .
The concept and method of PT-moments to detect entanglement was extended to define the realignment-moments Zhang-QIP21 as
@@
where $R(\rho)$ is the realigned matrix Rudolph-QIP4 ; Chen-QIC3 . As it is known that realignment criterion can detect specific bound entangled states, therefore the Hankel matrices based on moments of realigned matrix are expected to detect some PPT-entangled states. However, we should keep in mind that realignment criterion is unable to detect many other entangled states. Recently, the more general approach to this problem was adopted to extend PT-moments to positive-map moments (PM-moments) Wang-EPJ137 . If $\Lambda$ is a positive map but not completely positive then for a separable state $\rho_{AB}$, the matrix $\mathcal{I}\otimes\Lambda(\rho_{AB})$ must have all positive eigenvalues. This condition is both necessary and sufficient for detection of entanglement Horodecki-PLA223 . This means that if a given quantum state $\sigma_{AB}$ is entangled then there must exist a positive map, such that the matrix $\mathcal{I}\otimes\Lambda(\sigma_{AB})$ will have at least one negative eigenvalue. Therefore, it follows that the problem of detection of entanglement is to find the positive maps which are not completely positive. However, the issue is non-trivial because it is not easy to find the positive maps which are not completely positive. Even if we are able to find a specific positive map, it is not clear whether it will detect a given quantum state, because different positive maps may not detect entanglement of a state.
The moments of a positive map are defined as
@@
where $\Lambda$ is a positive map but not completely positive. This is a more general definition and PT-moments are a special case of this as transposition is a positive map but not completely positive. Hence, in general, if we have some positive maps which are not completely positive, then it is expected that moments of this maps should be able to detect some entangled states including PPT-entangled states. Once, we have the PM-moments, we can write the first two Hankel matrices
@@
We can write the two general criteria for entanglement detection as follows:
@@
Result 1: For a separable state $\rho_{AB}$, we have $S_{m}(q)\geq 0$ for $m=1,2$, etc., and violation means an entangled state for $S_{m}(q)<0$.
@@
This means that if $S_{m}(q)<0$, the state is entangled, however, it may be that $S_{m}(q)\geq 0$ for entangled states as well. We also note that in general the operation $\mathcal{I}\otimes\Lambda(\rho_{AB})$ may not be trace preserving, so if we want, we may normalize the mapped matrix by $(\mathcal{I}\otimes\Lambda(\rho_{AB}))/Tr(\mathcal{I}\otimes\Lambda(\rho_{AB}))$ with the consequence that $q_{1}=1$, otherwise $q_{1}\neq 1$. The normalization process is not really required as the positivity or negativity of a matrix does not depend on it. For a normalized matrix, the criterion proposed in Ref.Wang-EPJ137 is $q_{3}-q_{2}^{2}\geq 0$ for any positive map and for separable states. This criterion is called ’$q_{3}-\Lambda$’ criterion.
Result 2: For a separable state $\rho_{AB}$, the PM-moment $q_{3}$ satisfies the relation
$q_{3}\geq\,\alpha\,x^{3}+(1-\alpha\,x)^{3}\,,$ (20)
@@
where $x=[\alpha+\sqrt{\alpha((\alpha+1)q_{2}-1)}\,]/[\alpha(\alpha+1)]$ with $\alpha=\lfloor 1/q_{2}\rfloor$. This criterion is called optimal PM-moment criterion or ’$q_{3}-O\Lambda$’.
@@
In Ref.Wang-EPJ137 , the authors proposed ’$q_{3}-\Lambda$’ and ’$q_{3}-O\Lambda$’ as two criteria to detect entangled states. However, as they have also provided the example that both of these criteria are unable to detect the complete range of entanglement for the given quantum states. We provide details for this situation in subsection below, where we discuss entanglement detection for $3\otimes 3$ systems. We propose a modification in these criteria and suggest to use the Hankel matrices $S_{m}(q)$ for detection of entanglement such that $S_{m}(q)<0$ means entangled states with appropriate positive map for a given state. Because, as it is mentioned before, a different positive map may not be able to detect entanglement of a given state. For most of the cases, $S_{1}(q)$ and $S_{2}(q)$ would detect the complete range of entanglement. If needed, one can add more moments and check $S_{3}(q)$.
@@ V Discussion and Summary
We studied detection of bipartite and multipartite entanglement via moments of positive maps which are not completely positive. Such maps and their moments can detect entangled states. We proposed a slight modification in recently proposed criterion based on moments of positive maps to detect entanglement. Our proposal improved the earlier criterion to detect the complete range of entanglemnet for a specific family of states, where the earlier criterion was detecting most of entanglement but failed to do so for some range of the parameter. We provided many examples of bipartite quantum states which are completely detected by our proposal. We found that reduction map is equivalent to partial transpose map only for qubit-qubit system and for multiqubit system where map is applied to anyone of the qubit. It is obvious that neither PT-moments nor reduction map can detect PPT entangled states for three qubits. It remains to find the positive maps, which are capable to detect entanglement of PPT entangled states of three qubits. For higher dimensional systems reduction map is not equivalent to partial transpose map. There are still few issues with positive maps which are not completely positive. First, even if we find a positive map it is still not clear which entangled states it is capable of detecting. A different positive map may or may not detect the same entangled states. Second, for a given entangled state, it is not easy to find the positive maps which are not completely positive to detect the state. We aim to address these issues in our future work.

### v2

@@ I Introduction
Another issue in detection of entangled states in experiments is the quantum state tomography with large number of qubits. It is both hard and exhaustive to perform state tomography for many qubits. It was found that partial transpose moments (PT-moments) can quantify quantum correlations and they are more feasible than the state tomography Paris-LNP-2004 ; Tran-PRA92-2015 ; Tran-PRA94-2016 ; vanEnk-PRL108-2012 . In addition, it was found that PT-moments are related with randomized measurements and few PT-moments are enough to describe separable states Elben-PRA99 ; Brydges-Sc364 ; Elben-PRL125 ; Huang-NP16 . Based on moments of partial transpose matrix, two main criteria were worked out to detect bipartite entanglement Yu-PRL127 . First method was to use the lowest Hankel matrix leading to a criterion also defined in Elben-PRL125 . Second method is a necessary and sufficient condition which is based on some moments of partially transposed matrix. A set of inequalities, based on Hankel matrices, were also proposed to detect the entanglement of bipartite quantum states Neven-npjQI7 . The local randomized measurements are also proposed to detect multipartite entanglement Ketterer-PRL122 ; Ketterer-Q4 ; Ketterer-PRA106 ; Ohnemus-PRA107 ; Wyderka-PRL131 ; Zhang-arx and bound entanglement Imai-PRL126 . The idea of PT-moments was extended to realignment matrix for a quantum state. It was shown that the moments of a realigned matrix can detect some bound entanglement Zhang-QIP21 . Recently, an interesting method is proposed to detect entanglement in arbitrary dimensional bipartite quantum systems. The authors proposed to use moments of a Hermitian matrix $(\rho^{R})^{\dagger}\rho^{R}$, where $\rho^{R}$ is the well known realigned matrix for a given quantum state. It was shown that this method is capable of detecting entanglement for negative partial transpose states and also some bound entangled states Aggarwal-PRA109 ; Aggarwal-PRA108 . Recently, the relationship between principle minors and PT-moments is worked out where it was shown that all the PT-moments can be written as a simple function of principle minors Ali-QIP22 ; Zhang-AP534 ; Zhang-PRA108 . A more general approach was recently introduced to define the moments of positive maps which are not completely positive Wang-EPJ137 . Based on moments of positive maps, two criteria were proposed to detect entanglement. As transposition is a positive map which is not completely positive, therefore PT-moments criterion is a special case of this general criterion. Although, the criteria based on moments of positive map were able to detect entanglement for a specific family of quantum states, nevertheless there is some range of parameter where these criteria fail to detect entangled states Wang-EPJ137 .
@@
In this work, we propose a set of inequalities resulting from the condition that Hankel matrices are positive for separable states. If the determinant of second order Hankel matrix is negative then the given quantum state is entangled. This proposal is an extension of criterion for partial transpose moments Neven-npjQI7 to more general positive maps. As a result, we show that our suggestion is capable of detecting entanglement for several families of quantum states in various dimensions of Hilbert space. In addition, our proposal is capable to detect the complete range of entanglement for a specific family of states where the earlier criterion fails to do so Wang-EPJ137 . We find out that reduction map for qubits is equivalent to partial transpose map as the resulting matrices give the same set of eigenvalues for both maps. Consequently, the detection of entanglement is optimal for qubit-qubit system. We also extend the idea of moments of positive map to multipartite systems. In multiqubits systems, reduction map applied to anyone of the qubit is equivalent to partial transpose map. For GHZ and W states mixed with white noise, we find that all the resulting moments for reduction map are exactly the same as PT-moments.
@@ II Positive maps and their moments
Let $\mathcal{H}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ be the bipartite finite dimensional Hilbert space. A mixed bipartite quantum state $\varrho\in\mathcal{H}$ is a positive semidefinitive density matrix (no negative eigenvalues) and with unit trace. A density matrix is said to be separable (nonentangled) if it can be written as
@@
where $p_{i}\geq 0$ and $\sum_{i}\,p_{i}=1$. In general, it is not easy to verify for a given quantum state whether it is a convex combination of product states. To find out whether a given quantum state is entangled or not, is still an open question despite tremendous efforts in this area of research Horodecki-RMP81 ; Guehne-PR474 . A necessary condition for separability is to check the partial transpose of the density matrix. If $(\varrho^{AB})^{T_{B}}$ is negative (having atleast one negative eigenvalue) then state $\varrho^{AB}$ is entangled. This condition is necessary and sufficient for $2\otimes 2$ and $2\otimes 3$ quantum systems. However, for higher dimensions of Hilbert space, there are quantum states having positive partial transpose (PPT) nevertheless entangled. Recently, partial transpose criterion was reformulated via partial transpose moments (PT-moments) Elben-PRL125 . This was required to address the issue of state tomography with large number of qubits or higher dimensional systems. The PT-moments are defined Elben-PRL125 as
@@
where $k=1,2,\ldots,d$ with $d$ as dimension of Hilbert space $\mathcal{H}$. It was found that we do not have to compute all $d$ PT-moment, in fact only few PT-moments are sufficient to determine the separability properties of the state Elben-PRL125 . Let $B_{k}(P)$ are the $(k+1)\times(k+1)$ Hankel matrices containing moments $p_{k}$, such that we have
@@
A criterion was formulated based on the Hankel matrices that for separable states, Hankel matrices are positive, i.e., $B_{k}\geq 0$ Elben-PRL125 .
The two criteria for detection of entanglement worked out in Yu-PRL127 can be summarized as follows:
Result 1: For a separable state $\rho_{AB}$, we have $B_{m}(p)\geq 0$ for $m=1,2$, etc., and violation means an entangled state for $B_{m}(p)<0$. However, it may be that $B_{m}(p)\geq 0$ for entangled states as well. This criterion is called ’$p_{3}-$ criterion’.
Result 2: For a separable state $\rho_{AB}$, the PT-moment $p_{3}$ satisfies the relation
$p_{3}\geq\,\alpha\,y^{3}+(1-\alpha\,y)^{3}\,,$ (11)
where $y=[\alpha+\sqrt{\alpha((\alpha+1)p_{2}-1)}\,]/[\alpha(\alpha+1)]$ with $\alpha=\lfloor 1/p_{2}\rfloor$. This criterion is called optimal PT-moment criterion or ’$p_{3}$-OPPT criterion’Yu-PRL127 .
In our recent study, we have observed that for the majority of quantum states, only $B_{1}$ and $B_{2}$ are sufficient to determine the separability of a given state. In addition, it is obvious from the formulation of PT-moments that they can detect only the NPT region of quantum states and can not detect PPT-entangled states or even genuine entanglement in multipartite settings Ali-QIP22 . The above method of PT-moments to detect entanglement was extended to define the realignment-moments Zhang-QIP21 . Please note that the following method of entanglement detection is different from the recently proposed method in Aggarwal-PRA109 ; Aggarwal-PRA108 . The moments for realigned matrix are defined as
@@
where $R(\rho)$ is the realigned matrix Rudolph-QIP4 ; Chen-QIC3 . As it is known that realignment criterion can detect specific bound entangled states, therefore it is natural to expect that the Hankel matrices based on moments of realigned matrix may detect some PPT-entangled states. However, we should keep in mind that realignment criterion is unable to detect many other entangled states. In Zhang-QIP21 , the authors gave some examples to support their proposal and they also discussed the corresponding two relations proposed in Yu-PRL127 .
Recently, the more general approach to this problem was adopted to extend PT-moments to positive-map moments (PM-moments) Wang-EPJ137 . If $\Lambda$ is a positive map but not completely positive then for a separable state $\rho_{AB}$, the matrix $\mathcal{I}\otimes\Lambda(\rho_{AB})$ must have all positive eigenvalues. This condition is both necessary and sufficient for detection of entanglement Horodecki-PLA223 . This means that if a given quantum state $\sigma_{AB}$ is entangled then there must exist a positive map $\Lambda$, such that the matrix $\mathcal{I}\otimes\Lambda(\sigma_{AB})$ will have at least one negative eigenvalue. Transposition is one such positive map, which is not completely positive and hence can detect some entangled states. This means that PT-moments method is only an example for a larger set of positive maps which are not completely positive. Therefore, it follows that the problem of detection of entanglement is to find the positive maps which are not completely positive. This issue is non-trivial because it is not easy to find the positive maps which are not completely positive. Even if we are able to find such maps, it is not clear whether they will detect a given quantum state. Hence the challenge is to look for quantum state specific positive maps as other positive maps may not detect entanglement of a given state.
The moments of a positive map will be called PM-moments (positive map moments). These are defined as Wang-EPJ137
@@
where $\Lambda$ is a positive map but not completely positive. This is a more general definition and PT-moments can be regarded as a special case of this general formalism. Hence, in general, if we have some positive maps which are not completely positive, then it is expected that moments of this maps may be able to detect some entangled states including PPT-entangled states. Once, we have the PM-moments, we can write the first two Hankel matrices
@@
The authors in Wang-EPJ137 wrote the two general criteria for entanglement detection as below. We note that these two criteria are essentially simple extension of the above criteria by replacing PT with PM as follows:
@@
Result $\tilde{1}$: For a separable state $\rho_{AB}$, we have $S_{m}(q)\geq 0$ for $m=1,2$, etc., and violation means an entangled state for $S_{m}(q)<0$.
@@
This means that if $S_{m}(q)<0$, the state is entangled, however, it may be that $S_{m}(q)\geq 0$ for entangled states as well. The key difference with earlier criterion 1, mention above is the fact that in general the operation $\mathcal{I}\otimes\Lambda(\rho_{AB})$ may not be trace preserving (Transposition (including partial transpose) map preserves the trace of a matrix and we always have $p_{1}=1$). Therefore, we may normalize the mapped matrix by $(\mathcal{I}\otimes\Lambda(\rho_{AB}))/Tr(\mathcal{I}\otimes\Lambda(\rho_{AB}))$ with the consequence that $q_{1}=1$, otherwise $q_{1}\neq 1$. The normalization process is not really required as the positivity or negativity of a matrix does not depend on it. For a normalized matrix, the criterion proposed in Ref.Wang-EPJ137 is $q_{3}-q_{2}^{2}\geq 0$ for any positive map and for separable states. This criterion is called ’$q_{3}-\Lambda$’ criterion.
Result $\tilde{2}$: For a separable state $\rho_{AB}$, the PM-moment $q_{3}$ satisfies the relation
$q_{3}\geq\,\beta\,x^{3}+(1-\beta\,x)^{3}\,,$ (21)
@@
where $x=[\beta+\sqrt{\beta((\beta+1)q_{2}-1)}\,]/[\beta(\beta+1)]$ with $\beta=\lfloor 1/q_{2}\rfloor$. This criterion is called optimal PM-moment criterion or ’$q_{3}-O\Lambda$’. This is essentially the same result as above mentioned $p_{3}$-OPPT criterion for positive maps.
@@
Now we are in a position to present our main contribution in this article. The main motivation behind this work is the unability of earlier proposal to detect some entangled states. As an example, in Ref.Wang-EPJ137 , the authors proposed ’$q_{3}-\Lambda$’ and ’$q_{3}-O\Lambda$’ as two criteria to detect entangled states. However, in their work, the authors provided a family of entangled states and both of these criteria were unable to detect the complete range of entanglement for the given quantum states. We will provide details for this situation in subsection below, where we discuss entanglement detection for $3\otimes 3$ systems.
Proposal : Our proposal is in the similar lines to Ref. Neven-npjQI7 . We propose to use the Hankel matrices to detect entanglement or alternatively to check a set of inequalities resulting from determinant of the Hankel matrices $S_{m}(q)$ for detection of entanglement such that $S_{m}(q)<0$ means entangled states with appropriate positive map for a given state. Because, as it is mentioned before, a different positive map may not be able to detect entanglement of a given state. For most of the cases, $S_{1}(q)$ and $S_{2}(q)$ would detect the complete range of entanglement. If needed, one can add more moments and check $S_{3}(q)$.
Theorem : For a separable state the Hankel matrices $S_{m}(q)$ are positive and if any such matrix is negative then the state is entangled.
Proof: The proof of this theorem is very simple. For a separable state defined in Eq. (1) and for a positive map which is not completely positive, the evolved state $\varrho(t)=\mathcal{I}\otimes\Lambda(\varrho^{AB})=\sum_{i}^{n}\,p_{i}\,\varrho_{i}^{A}\otimes\Lambda(\varrho_{i}^{B})\,$, remains separable and therefore $S_{m}(q)>0$. Whereas for any other initial state $\rho^{AB}(t)=\mathcal{I}\otimes\Lambda(\rho^{AB})$, if $S_{m}(q)<0$ then state is gauranteed to be entangled as dictated by Horodecki’s theorem on separability Horodecki-PLA223 .
Lemma : For a normalized mapped matrix and for separable states, we have
$det\,\left(\begin{array}[]{ccc}1&q_{2}&q_{3}\\ q_{2}&q_{3}&q_{4}\\ q_{3}&q_{4}&q_{5}\end{array}\right)\,\geq 0\,.$ (25)
This Lemma is equivalent to say that the Hankel matrix $S_{2}(q)$ is positive for a separable state. An alternative procedure to verify the positivity of a matrix is via principal minor (PM), which is defined as the determinant of a sub-matrix obtained by removing some rows and columns from the main matrix. We denote the sub-matrices with symbols $S(i)$, $S(ij)$, and $S(ijk)$ and their determinants with symbols $[S(i)]$, $[S(ij)]$, and $[S(ijk)]$. A result in matrix theory says that if a matrix is positive semidefinite then all its principal minors are also positive semidefinite Horne . The positivity of Eq.25 implies that the following set of inequalities must be satisfied by separable states.
$q_{3}\geq 0\,,$
$q_{5}\geq 0\,,$
$q_{3}\geq q_{2}^{2}\,,$
$q_{3}\,q_{5}\geq q_{4}^{2}\,,$
$q_{5}\geq q_{3}^{2}\,,$
$q_{3}\,q_{5}-q_{4}^{2}-q_{2}^{2}\,q_{5}+2q_{2}\,q_{3}\,q_{4}-q_{3}^{3}\geq 0\,.$ (26)
The violation of any of this condition is equivalent to say that matrix $S_{2}(q)$ is negative and hence the respective quantum state is entangled. In next sections, we will provide several examples to show that this criterion is able to detect entangled states whereas earlier criteria fails to do so. Hence our proposal is an enhancement and improvement in this direction of entanglement detection.
@@ V Discussion and Summary
We studied detection of bipartite and multipartite entanglement via moments of positive maps. Positive maps which are not completely positive, can detect entangled states. However, this problem is not trivial because it is known that finding such maps is not an easy task. Even if we find such maps, it is not clear which entangled states they will optimally detect. It might be that some positive maps are specific to specific quantum states and/or some specific quantum states might have specific positive maps designed to detect them. We proposed a set of inequalities such that if any one of this inequality is violated then the given quantum state must be entangled. This proposal is an extension of partial transpose map inequalities and an improvement to the earlier criteria. We showed that our proposal is capable of detecting the complete range of entanglemnet for a specific family of states, whereas the earlier criteria failed to do so for some range of the parameter. We provided several examples of bipartite quantum states which are completely detected by our proposal. We found that reduction map is equivalent to partial transpose map only for qubit-qubit system and for multiqubit system where map is applied to anyone of the qubit. It is obvious that neither PT-moments nor reduction map can detect PPT entangled states for three qubits. It remains to find the positive maps, which are capable to detect entanglement of PPT entangled states of three qubits. For higher dimensional systems reduction map is not equivalent to partial transpose map. Recently, an interesting proposal investigated the detection of entanglement for bipartite quantum systems. The idea is to use moments of a Hermitian matrix $(\rho^{R})^{\dagger}\rho^{R}$, where $\rho^{R}$ is the well known realigned matrix Aggarwal-PRA109 ; Aggarwal-PRA108 . We have done some initial investigations whether this proposal can also detect the quantum states detected by our proposal. Our studies suggest that this more recent proposal is good to detect some specific quantum states, however it did not detected the quantum states which are detected by our proposal. It still remains to investigate whether using realignment criterion in this way or some other function of it, can be useful to detect every bound entangled state for bipartite as well as multipartite quantum state. We aim to address these issues in our future work.
Acknowledgments
The author is grateful to both reviewers for their constructive comments, useful suggestions and pointing to important references. The author appreciates the additional editorial evaluations.